- Is the following true or false? Justify your answer. (x XOR y) = xy +(x + y)
Answer #1
(x XOR y)'
= (xy'+x'y)'
= (xy')'(x'y)'
= (x'+y)(x'y)'
= (x'+y)(x+y')
= x'x+x'y'+yx+yy'
= 0+x'y'+yx+yy'
= x'y'+yx+yy'
= x'y'+yx+0
= x'y'+yx
= xy+x'y'
= xy + (x+y)'